Determine by direct integration the centroid of the area Shown. Express your answer in terms of a and b.

Y = kg x2 Fig. P5.41
Answer #1

heлд ч, ek, x1 Y = k 3 (x, y) = (a, b) b= kg a ² (x, y) = (a, b) ط اگه ا keb/ a3 Area A = S dy doc y = k 2 to y = kg x x=0 toCentroid == 1 y=ko to x 20 to a = = 1 ff in dy dx y = kg x² x=a ka x² x dy dx А ) 2=0 22 Put A aby , k = b k = b 12 a x a ba?y = kg 2² x=64 ) = * د ديا = Centroid y = 2 Ss y dx dy & x y zo to vob A y=0 x=(17) 2 Pul A= ab kg = 6 kbab 12 7 = 12 ab 1 TF 13 (22 638-16 - zadba = 12. - 3) - 12b4 (15-34) 12 a 6² 12 bx (15-14) ab 35 Centroid (57) = (0,60, 0,342